OpenStudy (anonymous):

can someone make r the subject please: V=(1/3)pi r^2 h

OpenStudy (anonymous):

v=(1/3)pir^2h. first can divide 1/3 to v which we get 3v=pir^2h. then we divide by pi and h which we get 3v/(pih)=r^2 from there we square root everything to both sides so we can solve for r, which we get the square root of 3v/(pih)=r.

OpenStudy (anonymous):

why when you divided by 3 its 3v?? i thought 3v is when you times it :S

OpenStudy (amistre64):

v ----- = v * 3/1 = 3v 1/3

OpenStudy (anonymous):

oh ok, thanks :)